OpenStudy (selena2345):

Evaluate The Expression x=-4,y=-8,z=3 -17+3-4

OpenStudy (anonymous):

wtf is that all it gives u

OpenStudy (anonymous):

That doesn't even make sense.. there are no letters in the actual equation..?

OpenStudy (anonymous):

ikr that y i said is that all it give u lol

OpenStudy (selena2345):

-17+z-x

OpenStudy (anonymous):

im in here but it says im not its weard

OpenStudy (anonymous):

the answer is -10

OpenStudy (anonymous):

i think it would be -18. Because : -17+3= -14 -14 + -4 = -18

OpenStudy (anonymous):

do uagree sophia

OpenStudy (anonymous):

i may not have been right let me try again .

OpenStudy (anonymous):

-18

OpenStudy (anonymous):

yeah i think it is -18.

OpenStudy (anonymous):

how did you get -10 ?

OpenStudy (anonymous):

the expression says -17+z-x key word minus and a negitive and a negitive makes a positive so the expression if u substitute would be -17+3+4 which equals -10 do u see where im comming from now

OpenStudy (anonymous):

Yes, but if you have a negative and you subtract something then you get a negative because its going left on a number line.

OpenStudy (anonymous):

http://www.chacha.com/askChaCha/-17%2B3%2B-4+%3D%3F <-- Check this out haha.

OpenStudy (anonymous):

noooo for example if ur doing 1- -1 when the is a minus sign and then a negitive no matter what the sign becomes a positive so it would become 1+1 u see what i mean now

OpenStudy (anonymous):

-17 + 3 -4 = 18 -17 - 4 = -21 +3 = 18 -17 + 3 = -14 - 4 = 18 PEMDAS

OpenStudy (anonymous):

so you got positive 18?

OpenStudy (anonymous):

@myfacewhen

OpenStudy (anonymous):

*rolls eyes* -18

OpenStudy (anonymous):

Ok, then i got the same answer as you.

OpenStudy (anonymous):

sorry @rduncan16 i disagree with you.

OpenStudy (anonymous):

math disagrees with him

OpenStudy (anonymous):

umm ur both dumb even though its pemdas u still have tofollow the rule that if there are to minus next to each other it becomes a positive how old r u i took this already ik what im talking about adn im 15 soo im already older than sophia so i know more than her and ps u did pemdas wrong it doesnt matter weather adition or subtraction goes first and again u still have to follow the ruld that if two negatives are by each other it becomes positive

OpenStudy (anonymous):

Woah, no need to be rude rduncan16, i never said you were dumb so no need to say i am. idc if you are 15 or not i said what i think the answer is and you said what you did she can see which answer she chooses, sheesh. And i have a 15 year old sister that agree's with me . Idc what you have to say anymore . I respected what you thought was the answer but you don't need to be calling me dumb and being disrespectful. wow what maturity,

OpenStudy (anonymous):

im very sorry its just that myfacewhen said math disagrees with me whichi infer is calling me stupid even though im best at math and that made me mad and im sorry i took it out on u

OpenStudy (anonymous):

Its okay, I think you are very smart, I just didn't agree with you on this problem and I do think that what myfacewhen said was uncalled for and i don't agree with him . Your not stupid, myfacewhen is just not very mature it seems.